# Susan Athey    Emil Palikot

### v4

@@ 3 Experimental Interventions and Theoretical Framework
Labor market transitions face two fundamental frictions: imperfect information about worker abilities and search costs in matching workers to opportunities. Workers address these through skill signaling (spence1978job) and professional networks (granovetter1973strength; ioannides2004job; pallais2016referential). For women entering male-dominated technology sectors, both mechanisms face gender-specific barriers. First, occupational stereotypes bias employer perceptions of female candidates’ technical abilities, creating information frictions even for qualified women (cheryan2013stereotypical; del2022more; murciano2022missing). Second, network structures systematically exclude women from informal professional connections that facilitate hiring, with women reporting smaller networks and facing both structural barriers and personal hesitation in networking behaviors that benefit men (greguletz2019women; forret2004networking). Third, women in male-dominated fields face psychological barriers that may reduce job search effectiveness. Women systematically under-report their technical abilities on resumes compared to similarly skilled men (murciano2022missing), use less self-promoting language in professional contexts (lerchenmueller2019gender), and may experience reduced confidence when confronting occupational stereotypes (spencer1999stereotype). We study the relative importance of these barriers by evaluating two interventions with distinct mechanisms: a mentoring program that addresses network gaps and builds confidence through connections to successful female tech professionals, and a scalable portfolio-building program designed to counter information frictions and stereotypes through credible skill signals.
@@
Poland’s technology sector employed 250,000 workers in 2021, with median monthly wages of USD 2,600, nearly triple the national median (gus2022). Despite strong labor demand driving 9% annual wage growth, women held only 30% of technology positions and comprised just 1.1% of the female labor force, compared to 5.2% of men (EC2021; fluff2021). This persistent underrepresentation mirrors patterns across developed economies where women hold approximately 30% of tech jobs and persists despite numerous initiatives supporting women in technology (unitednations2021).
@@ 3.1 Program Designs and Mechanisms
We partnered with Dare IT, a Polish social impact organization addressing gender barriers in technology employment, to evaluate two interventions targeting different constraints. Both programs serve women with baseline technical skills who have not yet secured technology jobs. The programs’ contrasting designs allow us to contrast network effects from signaling effects.
Mentoring: Network Expansion and Tacit Knowledge.
@@
Mentoring addresses gender-specific network barriers documented in prior research. Women in technology report smaller professional networks and face structural exclusions from informal networking opportunities as compared to their male colleagues (greguletz2019women). Female professionals may also hesitate to engage in networking behaviors that appear effective for men (forret2004networking). By connecting mentees to established female professionals, the program potentially overcomes both barriers and provides same-gender role models who can navigate gender-specific challenges in recruitment.
Our qualitative interviews with mentors reveal four operational channels of the Mentoring program: (i.) direct network expansion through mentor connections and introductions to hiring managers, (ii.) transmission of tacit knowledge about navigating tech sector recruitment as a woman, including addressing potential bias in interviews, (iii.) self-efficacy reinforcement to build confidence, and (iv.) technical skill development through code review. Mentors consistently emphasize networking as the primary mechanism, actively facilitating introductions and teaching mentees how to leverage professional connections effectively.
Challenges: Skill Signaling.
To overcome Mentoring’s capacity constraints, we designed Challenges as a scalable alternative targeting information frictions exacerbated by gender stereotypes. Program development began with structured interviews of hiring managers at twelve prominent Polish technology firms. These conversations revealed a systematic pattern: hiring managers perceived female applicants as lacking practical technical experience, even when formal credentials suggested otherwise. Managers emphasized that portfolio projects demonstrating hands-on capability could substantially alter hiring decisions, but expressed skepticism about generic online certifications. These perceptions align with documented gender bias in technical hiring: employers underestimate women’s programming abilities (murciano2022missing) and stereotypes about women’s technical competence create additional screening burdens for female candidates (cheryan2013stereotypical). A credible skill signal may be particularly valuable for women if it counters these negative stereotypes by providing concrete, verifiable evidence of technical proficiency.
@@
If network barriers constrain technology employment for some women, Mentoring should be especially effective for participants with weak existing professional connections: women in smaller cities and towns where local technology networks are sparse, those without family or friends working in technology, and those with longer gaps in employment who may have lost previous professional contacts. The network mechanism also predicts effects on observable network measures: in particular, we should observe increases in LinkedIn connection counts among Mentoring participants. Additionally, if experienced mentors’ networks generate value, treatment effects should vary with mentor characteristics. Mentors with longer technology sector tenure and those in managerial positions likely possess richer professional networks and better access to hiring decision-makers, potentially generating larger effects for their mentees.
@@ 4 Experimental Design and Data
We evaluate the Autumn 2021 cohort, for which recruitment opened in October 2021 and the program launched in January 2022. As in previous cohorts, applications substantially exceeded the number of available mentors. Even after Dare IT’s eligibility screening, which excluded applicants currently employed in technology and those lacking technical skills at a junior professional level, the applicant-to-mentor ratio remained 4:1.
@@
This oversubscription allowed us to implement a randomized evaluation while preserving Dare IT’s mentor-mentee matching process. We modified the standard selection procedure: instead of each mentor selecting one mentee (the historical practice), we asked mentors to select their two preferred candidates from the pool of eligible applicants. We then randomized within each pair, assigning one applicant to treatment (participation in Mentoring) and one to control (no participation). This paired randomization design controls for mentor selection preferences while generating exogenous variation in program access.
Of 160 mentors, 151 agreed to this modified selection process; 9 declined, because they did not want to spend effort selecting mentee who will not end up working with them. This yielded an initial sample of 302 applicants (151 treatment, 151 control) across 151 mentor pairs. However, 4 treatment-assigned participants secured technology jobs between randomization (November 2021) and program launch (January 2022). We excluded these individuals from analysis and allowed the initially-assigned control applicant to participate instead, maintaining program access for mentors who invested time in candidate selection. Our final experimental sample comprises 151 treatment participants and 147 control participants.
Appendix G provides robustness checks: (i.) excluding the 4 mentor pairs without matched controls, (ii.) including the 4 pre-program employment cases with outcomes set to 1, and (iii.) including the 9 non-participating mentors under the assumption their hypothetical mentees did not find technology jobs. Results are substantively unchanged across specifications.
@@
In addition to experimental subjects, we collected outcome data on 468 eligible applicants whom mentors did not select. These applicants provide a benchmark for assessing mentor selection effects independent of program participation. As shown in Figure 2, selected-but-control participants have substantially higher employment success than non-selected applicants, highlighting the importance of randomization for isolating treatment effects from selection effects.
@@
Challenges. We launched recruitment in December 2021, with the program scheduled to begin in February 2022. To generate initial interest, we guaranteed admission to the first 100 applicants. Registration was capped at 500 applicants, but high demand led to 525 applications before we closed the portal.
@@
Applicants self-certified that they (i.) were not currently employed in the technology sector, (ii.) possessed English language proficiency at B1 level or higher, and (iii.) had working knowledge of either Figma (User Experience Design track) or React (Front-end Development track). Unlike Mentoring, we did not manually verify technical skills prior to admission due to scalability considerations.
@@
We exclude the first 100 guaranteed-admission applicants from our analysis, as they were not subject to randomization. Among the remaining 425 applicants, we randomly assigned 200 to treatment (program access) and 225 to control (no access). Randomization employed stratification on four registration characteristics: track specialization (UX Design vs. Front-end Development), age category, residence type (Warsaw, other large city, or small town), and employment status (student, employed outside tech, or neither).
@@
Between randomization (December 2021) and program launch (February 2022), 17 treatment-assigned applicants secured technology employment. Consistent with our approach in Mentoring, we exclude these individuals from our primary analysis. Our final analytic sample includes 183 treatment participants and 225 control participants. Appendix G demonstrates that results are robust to including these 17 early-employment cases with outcomes coded as 1.
@@
Among the 183 treatment participants, 41 (22%) did not submit any program assignments, suggesting passive non-compliance. However, these individuals retained access to all program materials, weekly webinars with industry experts, and the Dare IT Slack community. Following intention-to-treat principles, we include them in our analysis regardless of active participation level.
We also collected outcome data on 160 applicants who submitted forms after enrollment closed. These late applicants were not randomized and do not contribute to our primary treatment effect estimates. However, they provide an additional comparison group for the off-policy evaluation in Section 6.
Appendix B reports covariate balance tests across treatment and control groups. Of 9 baseline characteristics, only social science degree shows a statistically significant imbalance at the 5% level, consistent with chance given multiple comparisons. Our main specifications control for all observed variables to improve precision.
@@ 4.1 Data
LinkedIn Data. Our primary outcome measure, pre-registered prior to data collection, is whether participants secured technology-sector employment following program completion. We collected employment data from participants’ LinkedIn profiles at two time points: four months post-program (May 2022 for Mentoring, June 2022 for Challenges) and twelve months post-program (May 2023 and June 2023, respectively).
@@
    Any position in a technology company—firms in software development, testing, and sales; data analytics; IT services; digital marketing; or online platforms—excluding finance, regulatory, legal, accounting, and HR roles.
@@
Self-employment, freelancing, and volunteer positions were not coded as new employment. Appendix D validates this classification by comparing average salaries for categories (1) and (2) using Glassdoor data; we find no statistically significant difference (p=0.9), supporting their equivalence.
Data collection and attrition.
We obtained complete LinkedIn data for all Mentoring participants at both measurement points. For Challenges, we collected data on all participants at four months and 384 of 408 participants (94%) at twelve months. Missing data resulted from participants changing LinkedIn profile URLs or privacy settings between recruitment and follow-up waves. Appendix G provides bounds and estimates under extreme attrition assumptions; main results remain statistically significant under all scenarios.
@@
Career changer 151 5.86 3.79 2 35 145 6.02 3.79 2 35
Years of tech experience 151 0.70 0.46 0 1 145 0.71 0.46 0 1
@@
Applicants both programs
@@ 5 Effectiveness of Dare IT Programs
Let each Dare IT applicant be characterized by $(X_{i},Z_{i},W_{i}^{M},W_{i}^{C},S_{i},Y_{i})$, where $i$ is the applicant’s index $i=1,..,\mathcal{I}$, $X_{i}\in\mathcal{X}$ is a matrix of characteristics of applicants that are observed by mentors and recorded in data, $\zeta_{i}\in\mathcal{Z}$ are characteristic of applicants that are not observed in the data (though some might be observed by mentors prior to treatment assignment), and $S_{i}\in\mathcal{S}$ is a random variable indicating whether the applicant has been selected by a mentor. [^7] To select mentees, mentors received information from the application form; the important part of the application that mentors observed, which we did not, are short introductory videos recorded by the applicants. Finally, $W_{i}^{P}\in\left\{0,1\right\}$ is the treatment assignment into program $P$ and $Y_{i}\in\mathcal{Y}$ is the observed outcome.
Let $Y_{i}(P)$ and $Y_{i}(0)$ be potential outcomes when applicant $i$ participates in program $P$ and when she does not; $P\in\left\{M,C\right\}$, where $M$ stands for Mentoring and $C$ for Challenges. We are primarily interested in estimating the magnitude of the average treatment effect, defined as $\tau^{P}\coloneqq E\left[Y_{i}(P)-Y_{i}(0)\right]$, and testing whether average treatment effects are higher than zero. We are also interested in comparing the treatment outcomes to one another, although our estimates of such contrasts are noisy due to the limited sample size in each experiment.
Selected Sample in Mentoring.
Dare IT asked mentors to choose two applicants they wanted to work with. Thus, selected applicants are not a random draw from the pool. Let $S_{i}$ be a random variable indicating whether a mentor selected the applicant $i$ or not; $S_{i}=1$ when the applicant has been selected by a mentor while $S_{i}=0$ otherwise. Only selected applicants were included in the randomization. Thus, we do not observe outcomes under treatment for the non-selected applicants. The average treatment effect of Mentoring that we estimate is defined as
$\tau^{M}\coloneqq E\left[Y_{i}(M)-Y_{i}(0)|S_{i}=1\right].$ (1)
Table 1 provides summary statistics of subjects in the experimental groups of the mentoring program and the non-selected applicants. There are notable differences between the mentoring control group (those selected by mentors ($S_{i}=1$) but not assigned to treatment ($W_{i}^{M}=0$) and the applicant group (where $S_{i}=0$ and thus, by construction, $W_{i}^{M}=0$), most notably in terms of the primary outcome of securing a tech job. In the Challenges program, participants were selected randomly from eligible applicants, where eligibility was determined based on observable characteristics.
5.1 Average Treatment Effects
In Figure 2, we display the proportions of participants employed in technology jobs over the span of thirteen months post-application to the program. This is illustrated across all experimental groups. [^8] We construct outcomes over time based on the new job’s start date. In some cases, we only observe the year the new job started rather than the specific month. For those observations, we assign December of that year as the starting month. across the experimental groups. In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT). We can see that all groups start at zero, with none of the applicants having a job in technology; over time, the share of subjects that have a tech job in each group increases, with the Mentoring group increasing right from the moment the program started, and achieving over 50% probability of having a job in tech after 8 months. In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points. This differential timing, with Mentoring effects emerging during the program while Challenges effects manifest several months post-completion, suggests distinct mechanisms. These trajectories are consistent with the distinction between immediate network-based job matching versus the creation of portfolio signals and gradual employer response to them.

### v5

@@ 3 Experimental Interventions and Theoretical Framework
Labor market transitions face two fundamental frictions: imperfect information about worker abilities and search costs in matching workers to opportunities. Workers can address these through skill signaling (Spence, 1978) and professional networks (Granovetter, 1973; Ioannides and Loury, 2004; Pallais and Sands, 2016). For women entering male-dominated technology sectors, both mechanisms face gender-specific barriers. First, occupational stereotypes bias employer perceptions of female candidates’ technical abilities, creating information frictions even for qualified women (Cheryan et al., 2013; Del Carpio and Guadalupe, 2022; Murciano-Goroff, 2022). Second, network structures systematically exclude women from informal professional connections that facilitate hiring, with women reporting smaller networks and facing both structural barriers and personal hesitation in networking behaviors that benefit men (Greguletz et al., 2019; Forret and Dougherty, 2004). Third, women in male-dominated fields face psychological barriers that may reduce job search effectiveness. Women systematically under-report their technical abilities on resumes compared to similarly skilled men (Murciano-Goroff, 2022), use less self-promoting language in professional contexts (Lerchenmueller et al., 2019), and may experience reduced confidence when confronting occupational stereotypes (Spencer et al., 1999). We study the relative importance of these barriers by evaluating two distinct interventions: a mentoring program that primarily focuses on network gaps and confidence building through connections to successful female tech professionals, and a scalable portfolio-building program designed to counter information frictions and stereotypes through credible skill signals.
@@
Poland’s technology sector employed 250,000 workers in 2021, with median monthly wages of USD 2,600, nearly triple the national median (GUS, 2022). Despite strong labor demand driving 9% annual wage growth, only 1.1% of women in the labor force held positions in the technology sector, compared to 5.2% of men (European Commission, 2021; Gawlowska-Bujok and Bujok, 2021). This persistent underrepresentation mirrors patterns across developed economies where women hold approximately 30% of tech jobs and persists despite numerous initiatives supporting women in technology (United Nations, 2021). As a high-income economy and a EU member state with gender gaps in technology employment comparable to Western European averages, Poland provides an appropriate setting for studying barriers facing women entering male-dominated sectors.
@@ 3.1 Program Designs and Mechanisms
We partnered with Dare IT, a Polish social impact organization addressing gender barriers in technology employment, to evaluate two interventions targeting different constraints. Both programs serve women with baseline technical skills who have not yet secured technology jobs.
Mentoring:
@@
Mentoring aims to alleviate gender-specific network barriers documented in prior research. Women in technology report smaller professional networks and face structural exclusions from informal networking opportunities as compared to their male colleagues (Greguletz et al., 2019). Female professionals may also hesitate to engage in networking behaviors that appear effective for men (Forret and Dougherty, 2004). By connecting mentees to established female professionals, the program potentially overcomes these barriers and provides same-gender role models who can navigate gender-specific challenges in recruitment.
Our qualitative interviews with mentors conducted prior to the experiment revealed four operational channels of the Mentoring program: (i.) direct network expansion through mentor connections and introductions to hiring managers, (ii.) transmission of tacit knowledge about navigating tech sector recruitment as a woman, including addressing potential bias in interviews, (iii.) self-efficacy reinforcement to build confidence, and (iv.) technical skill development through code review. Mentors consistently emphasize networking as the primary mechanism, actively facilitating introductions and teaching mentees how to leverage professional connections effectively.
Challenges:
To overcome Mentoring’s capacity constraints, we designed Challenges as a scalable alternative targeting information frictions exacerbated by gender stereotypes. Program development began with structured interviews of hiring managers at twelve prominent Polish technology firms. These conversations revealed a systematic pattern: hiring managers perceived female applicants as lacking practical technical experience, even when formal credentials suggested otherwise. Managers emphasized that portfolio projects demonstrating hands-on capability could substantially alter hiring decisions, but expressed skepticism about generic online certifications. These perceptions align with documented gender bias in technical hiring: employers underestimate women’s programming abilities (Murciano-Goroff, 2022) and stereotypes about women’s technical competence create additional screening burdens for female candidates (Cheryan et al., 2013). A credible skill signal may be particularly valuable for women if it counters these negative stereotypes by providing concrete, verifiable evidence of technical proficiency.
@@
If network barriers constrain technology employment for some women, Mentoring should be especially effective for participants with weak existing professional connections: women in smaller cities and towns where local technology networks are sparse, those without family or friends working in technology, and those with longer gaps in employment who may have lost previous professional contacts. The network mechanism also predicts expansion of professional networks, which we measure through LinkedIn connection counts. Additionally, if experienced mentors’ networks generate value, treatment effects should vary with mentor characteristics. Mentors with longer technology sector tenure and those in managerial positions likely possess richer professional networks and better access to hiring decision-makers, potentially generating larger effects for their mentees.
@@
Before proceeding, we note an important limitation: the two interventions do not isolate single mechanisms. Mentoring may include technical feedback alongside network expansion, while completing Challenges may affect confidence in addition to generating skill signals. Our empirical strategy therefore cannot definitively attribute effects to specific channels. Instead, we interpret differential treatment effects and heterogeneity patterns as evidence about the relative importance of network-based versus signaling-based approaches, recognizing that each program operates through multiple reinforcing channels.
@@ 4 Experimental Design and Data
We evaluate the cohort recruited in October 2021, with the program launching in January 2022. As in previous cohorts, applications substantially exceeded mentor availability. After Dare IT’s eligibility screening, which excluded applicants currently employed in technology and those lacking junior-level technical skills, the applicant-to-mentor ratio remained 4:1, enabling a randomized evaluation.
@@
We modified the standard selection procedure to generate exogenous variation in program access. Rather than each mentor selecting one mentee (the historical practice), we asked mentors to identify their two preferred candidates from eligible applicants. We then randomized within each pair in November 2021, assigning one applicant to treatment and one to control. This paired design controls for mentor selection preferences while ensuring balance on characteristics mentors value.
Of 160 mentors, 151 agreed to this modified process; 9 declined because they did not want to invest effort selecting candidates who might not become their mentees. This yielded 302 applicants (151 treatment, 151 control) across 151 mentor pairs. Between randomization and program launch, 4 treatment-assigned participants secured technology jobs and withdrew. To maintain program access for mentors who had invested time in selection, these slots were filled by the corresponding control-group applicants, yielding 151 treated participants and 147 controls. We assess sensitivity to this replacement procedure in Appendix G, estimating effects under alternative sample definitions; results are substantively unchanged.
@@
We also collected outcome data on 468 eligible applicants whom mentors did not select. These non-selected applicants provide a benchmark for assessing selection: control participants whom mentors chose exhibit substantially higher employment rates than non-selected applicants with similar eligibility, confirming that mentor selection predicts success and underscoring the importance of within-pair randomization for identification.
Challenges.
@@
We launched recruitment in December 2021, with the program beginning in February 2022. To generate initial interest, we guaranteed admission to the first 100 applicants. Registration was capped at 500, but high demand produced 525 applications before we closed enrollment.
@@
Applicants self-certified that they (i) were not currently employed in the technology sector, (ii) possessed English proficiency at B1 level or higher, and (iii) had advanced proficiency in either Figma (UX Design track) or React (Front-end Development track). Unlike Mentoring, we did not manually verify technical skills prior to admission, reflecting the program’s design for scalability.
@@
We exclude the 100 guaranteed-admission applicants from our experimental analysis. Among the remaining 425 applicants, we randomly assigned 200 to treatment (program access) and 225 to control (no access), stratifying on track specialization, age category, residence type (Warsaw, other large city, or small town), and employment status (student, employed outside tech, or neither).
@@
Between randomization and program launch, 17 treatment-assigned applicants secured technology employment and chose not to participate. Unlike Mentoring, where the paired design required replacing such individuals, here we simply exclude them from the primary analysis. Our final analytic sample includes 183 treatment and 225 control participants. Appendix G demonstrates robustness to including these 17 cases with outcomes coded as 1.
@@
Among treatment participants, 41 (22%) submitted no assignments, though they retained access to program materials, weekly webinars with industry experts, and the Dare IT Slack community. Following intention-to-treat principles, we include all treatment-assigned participants regardless of engagement level; our estimates therefore capture the effect of program access rather than active participation. Under standard LATE assumptions, effects for active participants would be proportionally larger.
We also collected outcome data on 160 applicants who registered after enrollment closed. These late applicants contribute to the off-policy evaluation in Section 6, which estimates treatment effects for populations beyond the experimental sample. Appendix B reports covariate balance; of 9 baseline characteristics, only social science degree shows imbalance at the 5% level, consistent with chance.
@@ 4.1 Data
LinkedIn Data.
Our primary outcome measure is whether participants secured technology-sector employment following program completion. We collected employment data from participants’ LinkedIn profiles until twelve months post-program (May 2023 and June 2023, respectively).
@@
    Any position in a technology company: firms in software development, testing, and sales; data analytics; IT services; digital marketing; or online platforms; excluding finance, regulatory, legal, accounting, and HR roles.
@@
Self-employment, freelancing, and volunteer positions were not coded as new employment. Appendix D validates this classification by comparing average salaries for categories (1) and (2) using Glassdoor data; we find no statistically significant difference between (1) and (2) (p=0.9).
We obtained complete LinkedIn data for all Mentoring participants. For Challenges, we collected data on 384 of 408 subjects (94%). Missing data resulted from participants changing LinkedIn profile URLs or privacy settings between recruitment and follow-up waves. Appendix G provides bounds and estimates under extreme attrition assumptions; main estimates remain statistically significant under all scenarios.
@@
Years of tech experience 151 5.86 3.79 2 35 145 6.02 3.79 2 35
Career changer 151 0.70 0.46 0 1 145 0.71 0.46 0 1
@@
Applicants both programs (Non-experimental comparison group)
@@ 5 Effectiveness of Dare IT Programs
Let each applicant $i=1,\ldots,N$ be characterized by observable characteristics $X_{i}\in\mathcal{X}$, unobservable characteristics $Z_{i}\in\mathcal{Z}$, treatment assignment $W_{i}^{P}\in\{0,1\}$ for program $P\in\{M,C\}$, and observed outcome $Y_{i}\in\mathcal{Y}$. Let $Y_{i}(P)$ denote the potential outcome under participation in program $P$, where $M$ denotes Mentoring and $C$ denotes Challenges, and let $Y_{i}(0)$ denote the potential outcome absent any program participation. For a subpopulation $\mathcal{G}\subseteq\{1,\ldots,N\}$, we define the average treatment effect as
$\tau^{P}_{\mathcal{G}}\coloneqq\mathbb{E}\left[Y_{i}(P)-Y_{i}(0)\mid i\in\mathcal{G}\right].$ (1)
Selection into Treatment.
The two programs employ different selection mechanisms, which affects the interpretation of our estimates. In the Challenges program, participants were randomly selected from all eligible applicants, where eligibility was determined by observable characteristics. Let $\mathcal{G}^{C}$ denote this eligible population. Random selection implies that our estimate of $\tau^{C}_{\mathcal{G}^{C}}$ identifies the average treatment effect for the eligible population.
In the Mentoring program, applicants were first screened by mentors who selected preferred mentees based on application materials, including short introductory videos, some of which we do not observe in the data. Let $\mathcal{G}^{S}\subseteq\{1,\ldots,N\}$ denote the subpopulation of applicants selected by mentors; only selected applicants entered the randomization, and thus we do not observe treated outcomes for non-selected applicants. This selection process implies that our estimand for the mentoring program is:
$\tau^{M}_{\mathcal{G}^{S}}\coloneqq\mathbb{E}\left[Y_{i}(M)-Y_{i}(0)\mid i\in\mathcal{G}^{S}\right].$ (2)
This quantity represents the average treatment effect for applicants whom mentors chose to work with, a population that may differ systematically from the broader applicant pool on both observable and unobservable dimensions. To the extent that mentors select applicants with higher potential returns to mentoring, $\tau^{M}(\mathcal{G}^{S})$ may exceed the average treatment effect that would obtain under random assignment from the full applicant pool.
5.1 Treatment Effects
Figure 2 displays technology employment rates over fourteen months following program application, separately for each program. [^8] These are simple unweighted averages. We drop from these plots subjects who reported only a year in which they started a new job. Appendix C presents the plots with these subjects included; the only group with noticeable difference are non-selected applicants to Mentoring. In the analysis, we imput month 14 for cases when the month is missing.
The left panel presents results for the Mentoring program, comparing three groups: treated applicants (selected by mentors and assigned to receive mentoring), control applicants (selected by mentors but assigned to the control group), and non-selected applicants (not chosen by any mentor and therefore excluded from randomization). All groups begin at zero as no applicants held technology jobs at baseline. Employment rates rise over time across all groups, but at markedly different rates. The treatment group exhibits rapid gains beginning immediately after program start, reaching 52% employment by month eight. The control group follows a similar but attenuated trajectory, stabilizing at 43%. Non-selected applicants show substantially lower employment throughout, reaching only 19% by the end of the observation period.